OpenStudy (anonymous):

Find the slope intercept form: x intercept 2 y intercept 2/3

hartnn (hartnn):

for slope intercept form y=mx+c, u need c=y-intercept = 2/3 and u need slope=m=? do u know how to find m ?

OpenStudy (anonymous):

change in y change in x ???

hartnn (hartnn):

to be precise m= - y-intercept / x-intercept can u find m now ?

OpenStudy (anonymous):

2/3 / 2 ?

hartnn (hartnn):

don't forget - *minus* sign

OpenStudy (anonymous):

Why is there a negative?

OpenStudy (anonymous):

-2/3 /2 Where does the negative come from?

hartnn (hartnn):

hmmm.... to get x-intercept, u put y=0 right ? in y=mx+c, if u put y=0, u get 0=m(x-int)+(y-int) so m= -(y-int)/(x-int)

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

Uhhmmm, No?

OpenStudy (anonymous):

then the intercept is -1/3?

hartnn (hartnn):

*slope* = m= -1/3

OpenStudy (anonymous):

I think I am making this harder than it is. You showed a new way of doing this. Thank you!

hartnn (hartnn):

but did u get the equation ?

OpenStudy (anonymous):

m=-(y-int)/(x-int)

hartnn (hartnn):

that is just m=-1/3. equation : y=mx+c y= (-1/3)x + 2/3 u can simplify this.

OpenStudy (anonymous):

-1x?

hartnn (hartnn):

by multiplying by 3 , u get 3y=-x+2 or x+3y=2 got it ?

OpenStudy (anonymous):

I guess not. Where did the 3 come from?

hartnn (hartnn):

because the fraction had denominator 3, to eliminate that, i multiply 3 on both sides....

OpenStudy (anonymous):

Ohhhh got it

hartnn (hartnn):

finally ! good :)

OpenStudy (anonymous):

Thanks for your patience.